OpenStudy (blackbird02):

Please help on graphing this function y= 3sin (2theta +pi/3) +5 More details below

OpenStudy (blackbird02):

How to graph these functions: a. y=\[3\sin (2\theta + \frac{ \pi }{ 3 })+5\] b. \[y = 3\sin (\frac{ \theta }{ 2}-\frac{ \pi }{ 6 }) - 4\] In this graph:

OpenStudy (blackbird02):

a.) Amplitude = 3 Period = π Phase Displacement = π/6 Vertical Translation = 5 b.) Amplitude = 3 Period = 4π Phase Displacement = -π/3 Vertical Translation = 4 Are these values correct and how do I graph them?

OpenStudy (mathmale):

I've noticed that your original function doesn't have an independent variable. Insert one, please.

OpenStudy (blackbird02):

@mathmale what do you mean? That's exactly how it is in the given/

OpenStudy (daniellelovee):

I don't know this exactly but this hopefully helps you a bit https://answers.yahoo.com/question/index?qid=20110508155922AAi6zs8

zepdrix (zepdrix):

Hey Bird :) You still around?

zepdrix (zepdrix):

a.) Amplitude = 3 Period = π Phase Displacement = π/6 Vertical Translation = 5 Information looks good! I would say that the phase displacement is -π/6 though, since it's going to be a horizontal shift to the left.

OpenStudy (blackbird02):

@zepdrix Hi, so how do I graph this? I'm just concerned on where the points will be in the x-axis.

zepdrix (zepdrix):

ok great you labeled your y-axis according to your amplitude. Let's leave the x-axis unlabeled for now, it will make things easier in the end.